Find the work done in moving an object along a vector r=3i+2j-5k, if the applied force is F=2i-j-k
1
Expert's answer
2019-10-14T12:49:01-0400

The work is equivalent to the dot product of the two vectors:

"W=rF=(3i+2j-5k)(2i-j-k)=\\=3*2+2*(-1)+(-5)*(-1)=9."